# Prompt Optimization via Retrieved Reasoning Assets and Multi-Agent Analysis

### v1

@@ 4.4. Baselines
We implement six baselines under three different categories: single-pass without retrieval (1, 2, and 3), retrieval-augmented without reasoning assets (4), and multi-agent prompt optimization (5 and 6). All methods generate an optimized prompt from the initial prompt, and the final response is produced from that optimized prompt by the same backbone LLM used for the method. Specifically, the six baseline methods are detailed below:
  1. (1)
    Direct Generation: The LLM rewrites the input prompt into an optimized one, without intermediate reasoning.
  2. (2)
    Chain-of-Thought (CoT) [2022] (Wei et al., 2022): The LLM performs step-by-step reasoning about how to improve the prompt, then outputs an optimized version.
  3. (3)
    Role Assignment [2023] (Shanahan et al., 2023): A crafted system role (e.g., “You are a meticulous assistant optimizing prompts for helpfulness, correctness, coherence, complexity, and verbosity.”) guides the LLM to output an optimized prompt in one pass.
  4. (4)
    Retrieval-Augmented Generation (RAG) [2020] (Lewis et al., 2020): M25 retrieves the top-$k=10$ training prompts with their responses and corresponding scores. The LLM references these exemplars to produce an optimized prompt.
  5. (5)
    MAD [2023] (Du et al., 2023): MAD (Multi-Agent Debate) framework organizes several LLM “debaters” that each propose an answer with a rationale, then engage in multi-round critiques of one another’s reasoning. A separate judge (or a consensus rule) evaluates the arguments at each round and selects the final answer, which measurably improves factuality and step-by-step reasoning on strategic-reasoning tasks.
  6. (6)
    MARS [2025] (Zhang et al., 2025): MARS is a hierarchical system of seven agents. A Planner designs the optimization trajectory, while a Teacher, Critic, Student Socratic dialogue iteratively refines the prompt. At each iteration, a Target module evaluates the candidates, logs the history, and the system finally outputs the best prompt.
Table 2. Main results on HelpSteer1 and HelpSteer2. Columns report the five HelpSteer metrics (Help, Corr, Coh, Comp, Verb) normalized to $[0,1]$ and their mean (Avg). Methods are grouped into (i) single-pass prompting, (ii) retrieval-augmented generation (no reasoning assets), and (iii) multi-agent frameworks. Best within each backbone is bold; second best is underlined.
Dataset Methods GPT-4o LLaMA-3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1 Direct Generation 0.3216 0.3866 0.7583 0.2951 0.4215 0.4366 0.2549 0.3247 0.7054 0.2702 0.4062 0.3927
Chain-of-Thought (CoT) 0.3223 0.3876 0.7595 0.2935 0.4192 0.4364 0.2359 0.3077 0.6906 0.2627 0.3967 0.3787
Role Assignment 0.3988 0.4679 0.8024 0.3427 0.5008 0.5025 0.2887 0.3516 0.7287 0.3164 0.4517 0.4274
RAG (sparse, k=10k{=}10) 0.3751 0.4402 0.7871 0.3116 0.4779 0.4784 0.2930 0.3594 0.7452 0.3007 0.4377 0.4272
MAD 0.3774 0.4764 0.7954 0.3210 0.5248 0.4990 0.3774 0.5049 0.8067 0.4084 0.6708 0.5531
MARS 0.4159 0.4876 0.7963 0.3293 0.5188 0.5096 0.3901 0.4591 0.7745 0.3512 0.5801 0.5110
HelpSteer2 Direct Generation 0.3616 0.4700 0.7723 0.3280 0.4913 0.4846 0.2616 0.3636 0.7078 0.2942 0.4421 0.4139
Chain-of-Thought (CoT) 0.2981 0.3958 0.7169 0.2888 0.4709 0.4341 0.1600 0.2545 0.6291 0.2421 0.3812 0.3334
Role Assignment 0.4400 0.5175 0.8221 0.4025 0.5992 0.5563 0.2555 0.3303 0.7050 0.3267 0.4441 0.4123
RAG (sparse, k=10k{=}10) 0.4903 0.5745 0.8642 0.4161 0.6567 0.6003 0.3990 0.4711 0.7989 0.3722 0.5814 0.5245
MAD 0.4167 0.5049 0.8067 0.4084 0.6708 0.5615 0.3971 0.4532 0.7898 0.3998 0.6439 0.5368
MARS 0.4791 0.5569 0.8268 0.4095 0.6234 0.5791 0.4296 0.5019 0.7994 0.3957 0.6181 0.5482
For all baselines, we adopt GPT-4o and LLaMA-3-8B-Instruct as the backbone models, ensuring consistency and fairness in comparison by aligning with our main experimental setup. Baseline experiments were conducted with the following settings for fairness and fairness of our framework: all LLM temperatures were set to 0 and all other settings were set to LLM defaults. For MAD, it was fixed as $agent=2$ and $round=3$, and MARS designed the planner’s steps in steps 343-4 and set the entire iteration to $5$.
Table 3. Ablation Studies on Retrieval Depth and Test Agent Combination. Varying the number of retrieved training examples ($k$) shows that $k=3$ (our default, denoted as MA-SAPO (k=3)) provides the best balance between contextual coverage and robustness. Collapsing the Analyzer and Refiner into a single agent yields competitive but less consistent results. Best is highlighted in bold; second best is underlined.
Dataset Methods GPT-4o LLaMA-3-8B
Help Corr Coh Comp Verb Avg Help Corr Coh Comp Verb Avg
HelpSteer1 $k=1$ 0.5182 0.6247 0.8543 0.4943 0.7349 0.6453 0.4045 0.4769 0.8268 0.4729 0.8446 0.6051
MA-SAPO (k=3) 0.5183 0.6260 0.8614 0.5013 0.7363 0.6486 0.4110 0.4868 0.8326 0.4720 0.8433 0.6091
HelpSteer2 $k=1$ 0.5058 0.5982 0.8458 0.5113 0.7402 0.6403 0.3910 0.4616 0.8215 0.4784 0.8420 0.5989
MA-SAPO (k=3) 0.5072 0.6038 0.8527 0.5244 0.7570 0.6490 0.4005 0.4754 0.8294 0.4833 0.8441 0.6065
5. Experimental Results and Analysis
5.1. Main Results
In our main experiments, we compare MA-SAPO with three families of baselines: (1) single-pass prompting methods including Direct Generation, Chain-of-Thought (CoT), and Role Assignment, (2) retrieval-augmented generation (RAG) without reasoning assets, and (3) multi-agent frameworks such as MAD and MARS. As outlined in Section 4.2, we evaluate on both HelpSteer1 and HelpSteer2 using the five annotated quality dimensions and their normalized average score as metrics. The results are shown in Table 2.
In detail, Single-pass methods provide the simplest form of optimization, but their weakness lies in performing edits in isolation. Direct Generation and CoT are essentially one-shot rewrites that lack diagnostic grounding, which makes their improvements brittle and often inconsistent across metrics. Role Assignment can guide the model to consider multiple aspects explicitly, yet it still suffers from over-reliance on handcrafted role descriptions and cannot adapt when those roles fail to cover task-specific subtleties. RAG improves upon these by injecting exemplars, but the optimization process remains unguided: retrieved examples are consumed without explanation of why they are useful, leaving the refinement process heuristic and prone to inconsistency. As a result, these methods either underperform on complex prompts or achieve higher scores only at the cost of verbosity and drift in task semantics.
Recent multi-agent frameworks such as MAD and MARS perform better by coordinating multiple roles in a structured workflow, yet they remain outcome-driven and heavy in cost [^5] A detailed cost and latency comparisons are provided in Section 5.4.. Their reliance on repeated debate or critique cycles incurs significant computational overhead, while still treating evaluation as a scalar target rather than a source of actionable reasoning. In contrast, our MA-SAPO consistently surpasses all baselines across datasets and backbones by explicitly converting evaluation signals into reusable reasoning assets. The Analyzer to Refiner pipeline applies only evidence-backed edits drawn from these assets, ensuring that optimizations are not only effective but also interpretable and controllable [^6] A detailed case study is provided in Section 5.5.. For instance, on HelpSteer1 with GPT-4o, MA-SAPO achieves 0.6486 average score compared to 0.5096 for MARS and 0.4784 for RAG. Similar margins hold across HelpSteer2 and with LLaMA-3-8B, confirming that structured reasoning, rather than trial-and-error search, is the main driver of performance. Overall, MA-SAPO combines accuracy and efficiency, making it a more practical and principled framework for prompt optimization than competitive baselines.
5.2. Ablation Study
5.2.1. Top-k Retrieval Variations
To examine how the number of retrieved training examples influences the performance of MA-SAPO, we vary the retrieval depth $k\in\{1,2,3,4\}$. Table 3 reports results for both GPT-4o and LLaMA-3-8B models across HelpSteer1 and HelpSteer2.
For both backbones, increasing $k$ from 1 to 3 generally improves overall performance. With $k=1$, the Analyzer is under-informed, as it can only compare the test prompt against a single exemplar, limiting the diversity of reasoning assets available. Moving to $k=2$ provides broader evidence and modest gains, while $k=3$ yields the best balance across all five metrics and achieves the highest average performance. At $k=4$, results remain competitive but show slight declines in some dimensions, suggesting that introducing too many exemplars adds noise and creates conflicting diagnostic signals.
These findings highlight a clear trade-off: smaller $k$ restricts contextual diversity, while larger $k$ risks diluting the relevance of retrieved assets. Overall, $k=3$ achieves the most consistent and robust performance across datasets and backbones, and we adopt it as the default retrieval setting in all subsequent experiments.
5.2.2. Test Agents Combination
To further assess the robustness of MA-SAPO, we conduct an ablation study where the two test-phase agents are merged into a single combined agent, referred to as Test Agents Combination. Table 3 summarizes the results for both GPT-4o and LLaMA-3-8B on HelpSteer1 and HelpSteer2.
We observe that the combined variant achieves competitive performance, occasionally surpassing the full MA-SAPO in a single metric. However, the average performance consistently falls behind the full framework. This suggests that while a unified agent can capture certain improvements, it lacks the structured diagnostic pipeline of MA-SAPO, where the Analyzer first produces evidence-grounded reports and the Refiner then executes targeted edits.
These findings demonstrate the necessity of separating the two roles in the test phase. The Analyzer to Refiner pipeline enforces a more interpretable and controllable workflow, ensuring that refinements are justified by retrieved reasoning assets rather than introduced implicitly by a single agent. As a result, the modular design of MA-SAPO is not only more reliable but also essential for achieving consistent gains across datasets and backbones.
5.3. Qualitative Analysis
@@
To complement the automatic benchmark results, we conducted a human evaluation designed to assess both the interpretability and controllability of MA-SAPO’s optimization process. This evaluation focused on two complementary hypotheses: (H1) that multi-agent reasoning produces higher-quality reasoning than a single-agent baseline, and (H2) that MA-SAPO preserves the semantic direction and intent of the original prompts during optimization.
@@
All evaluations were conducted by 14 independent human annotators (mean age = 29.1, SD = 3.4), who were either data scientists, AI engineers, or graduate students majoring in data science or related fields. Each annotator was familiar with language model evaluation and prompt optimization procedures. The evaluation employed a 5-point Likert scale for reasoning quality (H1) and a 4-point scale for directional consistency (H2). The following subsections describe the evaluation setup and results.
5.3.1. Evaluating Multi-Agent Reasoning Quality (H1)
@@
To assess whether MA-SAPO enhances reasoning quality, we randomly sampled 30 reasoning cases from the HelpSteer1 training set. Each case included two reasoning outputs: one generated by a single-agent baseline and one by MA-SAPO’s multi-agent reasoning pipeline, resulting in 60 reasoning outputs in total. These outputs were independently evaluated by 14 human annotators, yielding 840 individual evaluations (30 cases $\times$ 2 outputs $\times$ 14 annotators). Each output was rated along three qualitative dimensions:
@@ Results.
As summarized in Table 4, MA-SAPO outperformed the single-agent baseline across all three dimensions. The largest gains were observed in perceived usefulness and factual accuracy, both showing statistically significant improvements (*p* ¡ 0.05, paired $t$-test, $n=30$). These findings indicate that multi-agent reasoning yields more helpful, accurate, and consistent responses.
Table 4. Human evaluation of reasoning quality. MA-SAPO outperforms the single-agent baseline in usefulness, factual accuracy, and consistency. Scores are averaged on a 1–5 scale; asterisks denote significance levels from paired $t$-tests ($^{*}p<0.05$, $n=30$).
@@
$\Delta$ (Improvement) +0.25 +0.24 +0.21 +0.23
5.3.2. Evaluating Directional Consistency of Optimized Prompts (H2)
To evaluate whether MA-SAPO preserves the semantic intent of prompts during optimization, we randomly sampled 40 prompt pairs, each consisting of an original prompt ($A$) and its optimized version ($B$) generated by MA-SAPO. Each pair was independently rated by 14 human annotators, resulting in a total of 560 evaluations (40 pairs $\times$ 14 annotators). Annotators rated the directional consistency of each pair on a 4-point Likert scale (1 = completely changed, 2 = partially changed, 3 = mostly preserved, 4 = fully preserved), indicating how well the optimized prompt maintained the intent, goal, and domain of the original.
@@
Figure 3 visualizes the distribution of annotator ratings for directional consistency. Most scores are concentrated between 3 and 4, indicating that the optimized prompts largely preserved the original semantic intent. The mean rating was 3.36 with a standard deviation of 0.74 ($n=560$), demonstrating that MA-SAPO effectively enhances prompt clarity and structure while maintaining semantic stability. The relatively moderate standard deviation further suggests that annotators consistently judged the optimized prompts as semantically aligned with their originals.
@@
Figure 3. Distribution of directional consistency ratings (n=560). Ratings on a 4-point scale (1 = changed, 4 = preserved). The dashed line marks the mean (3.36) and the shaded band shows one standard deviation (SD = 0.74).
5.3.3. Summary of Human Evaluation

### v2

@@ Appendix B Baseline details
We compare our method against six baselines from three categories: single-pass prompting, retrieval-augmented generation, and multi-agent frameworks. For all methods, the model first produces an optimized prompt from the initial prompt, and the final response is then generated from the optimized prompt using the same backbone LLM.
B.1 Single-pass prompting
This category includes Direct Generation, Chain-of-Thought (CoT) (Wei et al., 2022), and Role Assignment (Shanahan et al., 2023). These methods optimize the prompt without retrieving external examples. Direct Generation rewrites the input prompt into an optimized prompt in a single pass. Chain-of-Thought (CoT) generates intermediate reasoning steps before producing the optimized prompt. Role Assignment prepends a predefined role or persona to the input prompt and generates the optimized prompt in a single pass.
B.2 Retrieval-augmented generation
This category includes RAG (Lewis et al., 2020). The model is provided with the top-$k=10$ retrieved training examples as references when generating the optimized prompt. No additional reasoning assets are used.
B.3 Multi-agent framework
This category includes MAD (Du et al., 2024) and MARS (Zhang et al., 2025). MAD uses multiple agents to propose candidate prompt revisions, exchange feedback over several rounds, and select the final optimized prompt. MARS uses a hierarchical multi-agent framework that first plans the optimization procedure and then iteratively refines the prompt.
For all baselines, we use GPT-4o and Llama3-8B-Instruct as the backbone models. We set the decoding temperature to 0 and keep all other generation settings at their default values. For MAD, we use $2$ agents and $3$ debate rounds. For MARS, the planner produces $3$ to $4$ planning steps, followed by $5$ refinement iterations.
Appendix C LLM-as-a-judge
To automatically evaluate the quality of responses generated from our HelpSteer-based prompt generation framework, we use ArmoRM-Llama3-8B-v0.1 as the LLM-as-a-judge. ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024a) is a Llama-3 8B-based reward model that departs from conventional scalar preference modeling by decomposing response quality into multiple interpretable reward objectives. Specifically, the model first learns attribute-level rewards from multi-dimensional absolute-rating data through an Absolute-Rating Multi-Objective Reward Model (ArmoRM), and then aggregates them into a final preference score using a Mixture-of-Experts (MoE) gating network conditioned on the prompt context.
This model is particularly suitable for our evaluation setting because it directly incorporates the five HelpSteer annotation dimensions (helpfulness, correctness, coherence, complexity, and verbosity) as internal reward objectives. Since the model is trained to predict these HelpSteer-related attributes as part of its multi-objective reward formulation, it is well aligned with our task of evaluating prompt-response pairs derived from a HelpSteer-based framework. In addition, its two-stage design separates attribute-level reward prediction from final scalar aggregation, which provides a more interpretable evaluation process than conventional single-score preference models.
Empirically, ArmoRM-Llama3-8B-v0.1 achieves an overall score of 89.0 on RewardBench, outperforming the corresponding Llama-3 8B Bradley–Terry reward model (83.6). These reported results support our use of ArmoRM-Llama3-8B-v0.1 not only as an open and efficient judge model, but also as a reward model whose internal objectives are structurally aligned with the HelpSteer evaluation dimensions.
Appendix D Hardware Specifications
Model Hardware Power (TDP) Access
Meta-Llama-3-8B-Instruct NVIDIA H100-80GB 700W Local GPU
GPT-4o OpenAI API
Table 5: Hardware specifications used for model inference.
Table 5 summarizes the hardware configurations used for model inference in our experiments. For the open-source model, inference with Meta-Llama-3-8B-Instruct was conducted on NVIDIA H100-80GB GPU. For the proprietary model, GPT-4o was accessed via the OpenAI API, and no local GPU resources were required.
Appendix E Per-backbone score distributions
Figures 5 and 6 decompose the aggregated KDE distributions from Figure 2 by backbone model, reporting score distributions for GPT-4o and Llama3-8B separately. In each figure, the five HelpSteer dimensions and their average are shown as individual subplots, comparing Direct Generation against MA-SAPO across both HelpSteer1 and HelpSteer2.
E.1 GPT-4o
Figure 5: KDE score distributions for GPT-4o. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution.
Under GPT-4o, MA-SAPO produces pronounced rightward shifts across all five dimensions. Helpfulness and complexity show the largest gains, with Direct Generation distributions concentrated around scores of 1.2–1.3 and MA-SAPO shifting the mass to 2.1–2.2. Verbosity exhibits the most dramatic absolute shift, with Direct Generation peaking near 1.9 and MA-SAPO reaching approximately 3.0. Correctness also improves substantially, with the distribution peak moving from roughly 1.8 to 2.5 while becoming more tightly concentrated. Coherence is the exception, as Direct Generation already achieves high scores near 3.0, leaving limited headroom for further improvement. Nevertheless, MA-SAPO still achieves a visible upward shift in the distribution mean. The Average subplot confirms that these per-metric gains compound into a consistent overall improvement across both HelpSteer1 and HelpSteer2.
E.2 Llama3-8B
Figure 6: KDE score distributions for Llama3-8B. Solid lines denote HelpSteer1 and dashed lines denote HelpSteer2. Vertical lines indicate the mean of each distribution.
Under Llama3-8B, the directional pattern is consistent with GPT-4o, though with notable differences in distribution shape. Correctness under Direct Generation exhibits a broad, flat distribution across low scores, whereas MA-SAPO produces a sharp, concentrated peak near 2.0, indicating not only higher scores but also more consistent output quality. Verbosity shows the most dramatic shift among all metrics, with the MA-SAPO distribution peak reaching approximately 3.4 compared to roughly 1.8 for Direct Generation. Helpfulness and complexity also shift rightward, though with somewhat smaller absolute gains than under GPT-4o. As with GPT-4o, coherence shows the most modest improvement, consistent with its already high baseline scores.
E.3 Cross-backbone comparison
Across both backbones, MA-SAPO consistently shifts score distributions toward higher values in all five dimensions. Verbosity, complexity, helpfulness, and correctness show the largest shifts, while coherence improves more modestly due to the high baseline scores already achieved by Direct Generation. One notable difference between backbones is that Llama3-8B shows a sharper distributional improvement in correctness, with Direct Generation producing a much flatter spread compared to GPT-4o. Overall, the per-backbone breakdown confirms that the improvements observed in the aggregated figure are consistent across model families.
Appendix F Qualitative analysis details
@@
This appendix provides the full details of the human evaluation summarized in Section 5.3. The evaluation focused on two complementary hypotheses: (H1) that multi-agent reasoning produces higher-quality reasoning than a single-agent baseline, and (H2) that MA-SAPO preserves the semantic direction and intent of the original prompts during optimization.
@@
All evaluations were conducted by 17 independent human annotators (mean age = 29.1, SD = 3.4), who were either data scientists, AI engineers, or graduate students majoring in data science or related fields. Each annotator was familiar with language model evaluation and prompt optimization procedures. The evaluation employed a 5-point Likert scale for reasoning quality (H1) and a 4-point scale for directional consistency (H2).
F.1 Evaluating multi-agent reasoning quality (H1)
@@
To assess whether MA-SAPO enhances reasoning quality, we randomly sampled 30 reasoning cases from the HelpSteer1 training set. Each case included two reasoning outputs: one generated by a single-agent baseline and one by MA-SAPO’s multi-agent reasoning pipeline, resulting in 60 reasoning outputs in total. These outputs were independently evaluated by 17 human annotators, yielding 1020 individual evaluations (30 cases $\times$ 2 outputs $\times$ 17 annotators). Each output was rated along three qualitative dimensions:
@@ Results.
As summarized in Table 6, MA-SAPO outperformed the single-agent baseline across all three dimensions. The largest gains were observed in perceived usefulness and factual accuracy, both showing statistically significant improvements ($p<0.05$, paired $t$-test, $n=30$). These findings indicate that multi-agent reasoning yields more helpful, accurate, and consistent responses.
@@
$\Delta$ (Improvement) +0.25 +0.24 +0.21 +0.24
Table 6: Human evaluation of reasoning quality. MA-SAPO outperforms the single-agent baseline in usefulness, factual accuracy, and consistency. Scores are averaged on a 1–5 scale; asterisks denote significance levels from paired $t$-tests ($^{*}p<0.05$, $n=30$).
F.2 Evaluating directional consistency of optimized prompts (H2)
To evaluate whether MA-SAPO preserves the semantic intent of prompts during optimization, we randomly sampled 40 prompt pairs, each consisting of an original prompt ($A$) and its optimized version ($B$) generated by MA-SAPO. Each pair was independently rated by 17 human annotators, resulting in a total of 680 evaluations (40 pairs $\times$ 17 annotators). Annotators rated the directional consistency of each pair on a 4-point Likert scale (1 = completely changed, 2 = partially changed, 3 = mostly preserved, 4 = fully preserved), indicating how well the optimized prompt maintained the intent, goal, and domain of the original.
@@
Figure 7 visualizes the distribution of annotator ratings for directional consistency. Most scores are concentrated between 3 and 4, indicating that the optimized prompts largely preserved the original semantic intent. The mean rating was 3.36 with a standard deviation of 0.74 ($n=680$), demonstrating that MA-SAPO effectively enhances prompt clarity and structure while maintaining semantic stability. The relatively moderate standard deviation further suggests that annotators consistently judged the optimized prompts as semantically aligned with their originals.
@@
Figure 7: Distribution of directional consistency ratings (n=680). Ratings on a 4-point scale (1 = changed, 4 = preserved). The dashed line marks the mean (3.36) and the shaded band shows one standard deviation (SD = 0.74).
F.3 Summary of human evaluation